# Youngjoon Lee1, Jinu Gong2, Joonhyuk Kang1 1School of Electrical Engineering, KAIST, South Korea

### v1

@@ Abstract
Kolmogorov–Arnold Networks (KAN) offer universal function approximation using univariate spline compositions without nonlinear activations. In this work, we integrate Error-Correcting Output Codes (ECOC) into the KAN framework to transform multi‐class classification into multiple binary tasks, improving robustness via Hamming‐distance decoding. Our proposed KAN with ECOC method outperforms vanilla KAN on a challenging blood cell classification dataset, achieving higher accuracy under diverse hyperparameter settings. Ablation studies further confirm that ECOC consistently enhances performance across FastKAN and FasterKAN variants. These results demonstrate that ECOC integration significantly boosts KAN generalizability in critical healthcare AI applications. To the best of our knowledge, this is the first integration of ECOC with KAN for enhancing multi-class medical image classification performance.
@@ I Introduction
The Artificial Intelligence (AI) has transformed healthcare through models that analyze medical data with unprecedented accuracy [1, 2]. Neural architectures from MLP [3] to Transformers [4] have revolutionized medical classification by employing their unique mathematical principles to effectively capture patterns in diverse healthcare datasets [5]. These technologies have advanced automated diagnosis in imaging and pathology, enhancing clinical decision support [6]. Despite their effectiveness, traditional neural networks lack transparency, hindering adoption in healthcare where decision interpretability is crucial [7]. Therefore, medical applications represent an ideal domain for interpretable architectures that maintain performance while providing clear mathematical reasoning.
Kolmogorov–Arnold Networks (KAN) [8, 9] have emerged as a promising alternative to MLP for machine learning tasks by leveraging the Kolmogorov-Arnold representation theorem [10]. Unlike conventional neural networks, KAN utilize univariate spline functions composed in a hierarchical structure to achieve universal function approximation without relying on nonlinear activation functions [11]. Furthermore, these networks offer enhanced interpretability through their explicit mathematical formulation while maintaining competitive performance across various tasks [12]. However, despite these advantages, KAN faces challenges in generalizability when applied to complex multi-class classification problems. Consequently, current KAN implementations often require meticulous hyperparameter tuning, which can lead to reduced generalization capability in real-world scenarios.
@@
Illustration of our proposed KAN with ECOC framework is presented in Figure 1, demonstrating how we combine the structural advantages of KAN with the error-correcting properties of ECOC. The framework transforms the original multi-class problem into a set of binary classification tasks, where each binary classifier corresponds to a bit position in the codeword matrix. During training, KAN learns to predict individual bits of the codeword rather than directly predicting class labels. At inference time, the binary outputs from all KAN form a predicted codeword which is decoded to the nearest valid class codeword according to Hamming distance. Additionally, this approach provides robustness against individual classifier errors while maintaining the interpretability benefits of KAN.
@@
  • Additionally, ablation studies demonstrate that ECOC improves performance across KAN variants—including FastKAN [18] and FasterKAN [19]—confirming its versatility.
@@ II System Model
In this section, we present KAN for universal function approximation and ECOC for error-correcting multi-class decomposition. Our framework integrates these approaches to enhance classification generalizability in medical imaging applications.
@@ II-A Kolmogorov-Arnold Networks
where $z_{j}^{(l)}$ is the $j$-th output of layer $l$, $d_{l-1}$ is the previous layer’s dimensionality, and $g_{ij}^{(l)}$ are learnable univariate functions implemented as B-spline curves defined by control points.
@@
KAN training proceeds by minimizing the cross-entropy loss function:
@@ II-B Error-Correcting Output Codes
ECOC transforms multi-class classification into multiple binary classification tasks by encoding class labels as binary codewords with sufficient Hamming distance between classes. Specifically, for a $k$-class problem, ECOC defines a coding matrix $M\in\{-1,1\}^{k\times b}$, where $b$ is the codeword length, typically set to $b=2k$ to ensure sufficient code separation for optimal error correction.
The ECOC framework consists of two principal components:
@@
    $\hat{y}=\arg\min_{i\in\{1,\ldots,k\}}d_{H}(M_{i},\hat{C}),$ (6)
    where the Hamming distance between two binary codewords is defined as:
    $d_{H}(M_{i},\hat{C})=\sum_{j=1}^{b}\mathds{1}(c_{ij}\neq\hat{c}_{j}).$ (7)
For binary classification tasks within the ECOC framework, we employ the binary cross-entropy loss function:
@@ II-C Integration Framework
Our framework integrates KAN with ECOC to address generalizability challenges in multi-class classification. The integration methodically decomposes the original multi-class problem into binary classification tasks, trains separate KAN models for each binary task, and combines their outputs during inference.
As illustrated in Figure 1, the framework consists of the following components:
@@
    Coding Matrix Construction: We use the previously defined coding strategy to generate matrix $M\in\{-1,1\}^{k\times 2k}$.
@@
This integration simplifies classification tasks, enhances robustness through ECOC error-correction, and reduces hyperparameter sensitivity, collectively improving model generalization.
@@ III-A Experiment Setting
We evaluated our KAN with ECOC framework using the blood cell classification dataset [20], which contains microscopic images of eight different blood cell types. Experiments were conducted across different hyperparameter settings, including spline orders ($s$), grid sizes, and hidden layer configurations. All models were trained for 100 epochs using Adam optimizer with a learning rate of $0.001$, with experiments repeated across six different random seeds. Note that, we utilized EfficientKAN for efficient computation. Our implementation is available in our open source repository [^1] https://github.com/yjlee22/kan-ecoc.
@@ III-B1 Impact of ECOC
To investigate the impact of ECOC integration on KAN performance, we analyzed the results in Table I comparing vanilla KAN with our proposed approach. Our KAN-ECOC framework consistently outperforms the vanilla KAN across all metrics, demonstrating a clear performance advantage. The accuracy and recall show identical improvements from 0.8309 to 0.8519, while precision exhibits the most substantial gain from 0.8372 to 0.8690. Furthermore, the F1 score increases by 0.0261 points from 0.8320 to 0.8581, confirming the overall effectiveness of our approach. Additionally, the lower standard deviations observed across all metrics for KAN with ECOC indicate enhanced model stability and reliability, suggesting better generalization capabilities for medical image classification tasks.
@@ III-B2 Impact of Hyperparameter Configuration
To check the effect of hyperparameter variations on model performance, our experiments reveal that KAN with ECOC consistently outperforms vanilla KAN across all tested hyperparameter combinations. As shown in Fig. 2, the performance gap is maintained regardless of grid size (3, 5, or 10), with the most significant improvements observed in deeper network architectures ([5,5] and [5,5,5]). While both methods benefit from increased spline complexity (s=2, s=3), our ECOC integration provides additional gains at every spline order level. Notably, ECOC’s performance advantage remains consistent across all grid resolutions, demonstrating that our approach enhances generalization independently of this parameter. Additionally, the smaller error bars for KAN with ECOC indicate reduced variance and greater reliability in classification tasks, complementing the improved accuracy.
Figure 3: F1 scores (mean $\pm$ std) comparing vanilla and ECOC-integrated versions of FastKAN and FasterKAN across different grid sizes.
@@ IV Conclusion
In this paper, we introduce a novel framework integrating KAN with ECOC to enhance multi-class classification performance in medical imaging applications. The performance gains remain stable across various hyperparameter settings, demonstrating our framework’s robustness to grid sizes, spline orders, and network architectures. Additionally, ablation studies with FastKAN and FasterKAN variants confirm that ECOC integration provides consistent benefits regardless of the underlying KAN implementation. These comprehensive results validate our approach’s effectiveness for healthcare AI applications where both performance and stability are essential requirements.

### v2

@@ Abstract
Kolmogorov–Arnold Networks (KAN) offer universal function approximation using univariate spline compositions without nonlinear activations. In this work, we integrate Error-Correcting Output Codes (ECOC) into the KAN framework to transform multi‐class classification into multiple binary tasks, improving robustness via Hamming distance decoding. Our proposed KAN with ECOC framework outperforms vanilla KAN on a challenging blood cell classification dataset, achieving higher accuracy across diverse hyperparameter settings. Ablation studies further confirm that ECOC consistently enhances performance across FastKAN and FasterKAN variants. These results demonstrate that ECOC integration significantly boosts KAN generalizability in critical healthcare AI applications. To the best of our knowledge, this is the first work of ECOC with KAN for enhancing multi-class medical image classification performance.
@@ I Introduction
In recent years, deep learning has transformed healthcare through models that analyze medical data with unprecedented accuracy [1, 2]. Neural architectures from MLP [3] to Transformers [4] have revolutionized medical classification by employing their unique mathematical principles to effectively capture patterns in diverse healthcare datasets [5]. These AI models have advanced automated diagnosis in imaging and pathology, enhancing clinical decision [6]. Despite their effectiveness, traditional neural networks lack transparency, hindering adoption in healthcare where decision interpretability is crucial [7]. Therefore, medical applications represent an ideal domain for interpretable architectures that maintain performance while providing clear mathematical reasoning.
Kolmogorov–Arnold Networks (KAN) [8, 9] have emerged as a promising alternative to MLP for deep learning tasks by leveraging the Kolmogorov-Arnold representation theorem [10]. Unlike conventional neural networks, KAN utilize univariate spline functions composed in a hierarchical structure to achieve universal function approximation without relying on nonlinear activation functions [11]. Furthermore, these networks offer enhanced interpretability through their explicit mathematical formulation while maintaining competitive performance across various tasks [12]. However, despite these advantages, KAN faces challenges in generalizability when applied to complex multi-class classification problems. Consequently, current KAN implementations often require meticulous hyperparameter tuning, which can lead to reduced generalization capability in real-world scenarios.
@@
Illustration of our proposed KAN with ECOC framework is presented in Fig. 1, demonstrating how we combine the structural advantages of KAN with the error-correcting properties of ECOC. The framework transforms the original multi-class problem into a set of binary classification tasks, where each binary classifier corresponds to a bit position in the codeword matrix. During training, KAN learns to predict individual bits of the codeword rather than directly predicting class labels. At inference time, the binary outputs from all KAN form a predicted codeword which is decoded to the nearest valid class codeword according to Hamming distance. Additionally, this approach provides robustness against individual classifier errors while maintaining the interpretability benefits of KAN.
@@
  • Additionally, ablation studies demonstrate that ECOC improves performance across KAN variants—including FastKAN [18] and FasterKAN [19].
@@ II System Model
In this section, we present KAN for universal function approximation and ECOC for error-correcting multi-class decomposition. Our framework integrates these approaches to enhance general classification performance and robustness in challenging medical imaging applications.
@@ II-A Kolmogorov-Arnold Networks
where $z_{j}^{(l)}$ is the $j$-th output of layer $l$, $d_{l-1}$ is the previous layer’s dimensionality, and $g_{ij}^{(l)}$ are learnable univariate functions implemented as B-spline curves.
@@
The KAN training proceeds by minimizing the cross-entropy loss function:
@@ II-B Error-Correcting Output Codes
The ECOC transforms multi-class classification into multiple binary classification tasks by encoding class labels as binary codewords with sufficient Hamming distance between classes. In detail, for a $k$-class problem, ECOC defines a coding matrix $M\in\{-1,1\}^{k\times b}$, where $b$ is the codeword length, typically set to $b=2k$ to ensure sufficient code separation for optimal error correction. The ECOC consists of two principal components:
@@
[…]
    where the Hamming distance between two binary codewords is defined as $d_{H}(M_{i},\hat{C})=\sum_{j=1}^{b}\mathds{1}(c_{ij}\neq\hat{c}_{j}).$
Subsequently, for binary classification tasks within the ECOC method, we employ the binary cross-entropy loss function:
@@ II-C Integration Framework
Our framework integrates KAN with ECOC to address generalizability challenges in medical multi-class classification task. The integration methodically decomposes the original multi-class problem into binary classification tasks, trains separate KAN models for each binary task, and combines their outputs during inference.
As illustrated in Fig. 1, the framework consists of the following components:
@@
    Coding Matrix Construction: The proposed method uses the previously defined coding strategy to generate matrix $M\in\{-1,1\}^{k\times 2k}$.
@@
This integration enhances robustness through ECOC error-correction, and reduces hyperparameter sensitivity, collectively improving model generalization.
@@ III-A Experiment Setting
We evaluated our KAN with ECOC framework using the blood cell classification dataset [20], which contains microscopic images of eight different blood cell types. Experiments were conducted across different hyperparameter settings, including spline orders ($s$), grid sizes, and hidden layer configurations. All models were trained for 100 epochs using Adam optimizer with a learning rate of $0.001$, with experiments repeated across six different random seeds. Note that, we utilized EfficientKAN for computational efficiency. Our implementation is available in our open source repository [^1] https://github.com/yjlee22/kan-ecoc.
@@ III-B1 Impact of ECOC
To investigate the impact of ECOC integration on KAN performance, we analyze the results in Table I comparing vanilla KAN with our proposed approach. Our KAN-ECOC framework consistently outperforms the vanilla KAN across all metrics, demonstrating a clear performance advantage. The accuracy and recall show identical improvements from 0.8309 to 0.8519, while precision exhibits the most substantial gain from 0.8372 to 0.8690. Furthermore, the F1 score increases by 0.0261 points from 0.8320 to 0.8581, confirming the overall effectiveness of our approach. The results indicate that ECOC integration strengthens KAN with improved generalization.
@@ III-B2 Impact of Hyperparameter Configuration
To investigate how hyperparameter variations influence model performance, we conducted experiments across diverse configurations. As shown in Fig. 2, the performance gap is maintained regardless of grid size (3, 5, or 10), with the most significant improvements appearing in deeper networks ($[5,5]$ and $[5,5,5]$). While both methods benefit from increased spline complexity ($s=2$, $s=3$), our ECOC integration provides additional gains at every spline order level. Notably, ECOC’s performance advantage remains consistent across all grid resolutions, demonstrating that our approach enhances generalization independently of this parameter. Thus, KAN-ECOC provides more stable performance across varying hyperparameter settings.
Figure 3: Comparison of F1 scores (mean $\pm$ std) with vanilla and ECOC-integrated versions of FastKAN and FasterKAN across different grid sizes.
@@ IV Conclusion
In this work, we introduce a novel framework integrating KAN with ECOC to enhance multi-class classification performance in medical imaging applications. The performance gains remain stable across various hyperparameter settings, demonstrating our framework’s robustness to grid sizes, spline orders, and network architectures. Additionally, ablation studies with FastKAN and FasterKAN variants confirm that ECOC integration provides consistent benefits regardless of the KAN implementation. As future work, we plan to extend KAN–ECOC to federated learning, enabling privacy-preserving solutions in distributed healthcare settings.